OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation 7x – 5y = 28? Choose all answers that are correct. A. (−6, −14) B. (−1, −7) C. (4, 10) D. (7, 9)

OpenStudy (jojokiw3):

A is right. B is right.

OpenStudy (anonymous):

Thank you <33

OpenStudy (anonymous):

a and b both have each as negative numbers though

OpenStudy (jojokiw3):

Yeah, but look at this. 1 x 1 = 1. -1 x 1 = -1 1 x -1 = -1 -1 x -1 = 1